# PreCurious: How Innocent Pre-Trained Language Models Turn into Privacy Traps

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690279isbn: 979-8-4007-0636-3/24/10ccs: Security and privacyccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
We demonstrate that our attack can successfully amplify various privacy risks. Figure 1 illustrates the increased privacy vulnerability of MIA by our methods on both Head-FT and Adaptor-FT. More generally, for MIA, we compare PreCurious with benign GPT-2 (Radford et al., 2019) on the same black-box attack and demonstrate that by manipulating the pre-trained model, the true-positive-rate (TPR) at a false-positive-rate (FPR) of $0.01\%$ on Enron (Klimt and Yang, 2004), PubMed (Cohan et al., 2018) and PTB (Marcus et al., 1993) datasets is boosted by 8×,131×8$\times$,131$\times$8 × , 131 × and 36×36$\times$36 ×, respectively. For untargeted data extraction attack, we increase the times for a less duplicated sub-sequence shown in the pool of filtered generations by around 10×10$\times$10 ×. For targeted data extraction attack on Enron dataset, fine-tuning over benign model initialization cannot expose any secrets when fine-tuning with a strong DP level ($\epsilon=0.05$) while PreCurious can extract 3 target email addresses with valid exposure values. As advocated by previous work (Tramèr et al., 2022), we also audit the stealthiness of PreCurious and propose a mitigation method.
@@ 2.1. Parameter-Efficient Fine-tuning (PEFT)
One line of selective PEFT selects a portion of parameters in $\theta_{\text{pre}}$ as $\Phi$, such as Head-FT with a few top layers (Donahue et al., 2014) and Bitfit-FT with the bias terms of the model (Zaken et al., 2021). The other line of PEFT introduces new randomly initialized modules as $\Phi$ as plug-in for $\theta_{\text{pre}}$. For example, additive method Adapter-FT (Houlsby et al., 2019) inserts small and trainable fully connected networks $\Phi$ after transformer sub-layers in $\theta_{\text{pre}}$. The reparameterization-based method LoRA-FT (Hu et al., 2021) employs a low-rank matrix decomposition to parameterize the weight updates, and $\Phi$ indicates the parameters for the low-rank matrices.
@@ 2.3. Success Metrics
Now we introduce concrete attack effectiveness metrics for different attacks and propose stealthiness metrics for the victim to audit the pretrained model.
@@ 4.2.1. Performance Comparison
In addition, we observe the advantage introduced by model initialization in Table 3 by comparing Benign with Basic and the more aggressive Accelerated. We set $E_{\text{pre}}=1$ as a safe choice for the accelerated version on all datasets. There is a clear trend that the Val-PPL is decreasing and the privacy risk is increasing from Benign to Basic to Accelerated. The Accelerated is indeed a more aggressive strategy that pushes the starting point to the memorization-only stage.
@@ 4.2.4. Effectiveness under Defense
Figure 8. Duplication statistics and MIA effectiveness with Full-Ref under deduplication defense on $D_{\text{ft}}$ for PTB. The stealthiness metric $S_{\text{gap}}=\Delta_{\text{ppl}}$ is linearly scaled for clear visualization. We randomly subsample non-membership samples for keeping the same as the deduplicated $D_{\text{ft}}$ in MIA evaluation.
@@
Figure 9. MIA effectiveness on PTB dataset with $D_{\text{aux}}^{\text{dedup}}$ and $D_{\text{aux}}^{\text{dup}}$ as auxiliary data for training $\theta_{\text{ref}}$ in Benign or $\theta_{\text{pre/ref}}$ in PreCurious. The duplication level is increasing from left to right, with $|D_{\text{dedup/dup}}|/|D_{\text{ft}}|=1$ as the default $D_{\text{aux}}$ w/o deduplication.
Under the representative defense strategy of weight decay, we show in Table 5 that PreCurious is robust on privacy risk amplification, supported by more results w.r.t. various coefficient factors in the Appendix Table 8.
Under the strict defense of DP fine-tuning (Yu et al., 2021; Li et al., 2021), we show in Table 6 that PreCurious model increases the AUC compared to the Benign model but has a smaller TPR@0.01FPR and better utility due to the warming-up. The overall risk compared to non-DP fine-tuning in Table 1 is significantly mitigated by DP.
@@ 4.2.6. Breaking-up the trade-off
In addition, we address the privacy-utility trade-off issue in Table 6 with the lagging strategy as shown in Figure 10. Even when the worst-case privacy is bounded by a strict DP, we show that $\epsilon=0.05$ is still not a perfect protection. This success is due to more iterations for achieving the same utility, and also because the larger gradient norm derived from PreCurious-lagging fully exploits the per-sample sensitivity to reflect the influence of each sample.
@@ Appendix A Experimental Setup and More Results
Our research experiments were conducted on a system running Ubuntu 20.04.6, which was equipped with 8 NVIDIA Quadro RTX 8000 GPUs, each with 48,601 MiB of memory. The source code and other artifacts have been made available  [^2] https://github.com/Emory-AIMS/PreCurious.

### v3

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690279isbn: 979-8-4007-0636-3/24/10ccs: Security and privacyccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
We demonstrate that our attack can successfully amplify various privacy risks. Figure 1 illustrates the increased privacy vulnerability of MIA by our methods on both Head-FT and Adaptor-FT. More generally, for MIA, we compare PreCurious with benign GPT-2 (Radford et al., 2019) on the same black-box attack and demonstrate that by manipulating the pre-trained model, the true-positive-rate (TPR) at a false-positive-rate (FPR) of $0.01\%$ on Enron (Klimt and Yang, 2004), PubMed (Cohan et al., 2018) and PTB (Marcus et al., 1993) datasets is boosted by 8×,131×8$\times$,131$\times$8 × , 131 × and 36×36$\times$36 ×, respectively. For untargeted data extraction attack, we increase the times for a less duplicated sub-sequence shown in the pool of filtered generations by around 10×10$\times$10 ×. For targeted data extraction attack on Enron dataset, fine-tuning over benign model initialization cannot expose any secrets when fine-tuning with a strong DP level ($\epsilon=0.05$) while PreCurious can extract 3 target email addresses with valid exposure values. As advocated by previous work (Tramèr et al., 2022), we also audit the stealthiness of PreCurious and propose a mitigation method to make it more stealthy.
@@ 2.1. Parameter-Efficient Fine-tuning (PEFT)
One line of selective PEFT selects a portion of parameters in $\theta_{\text{pre}}$ as $\Phi$, such as Head-FT with a few top layers (Donahue et al., 2014) and Bitfit-FT with the bias terms of the model (Zaken et al., 2021). The other line of PEFT introduces new randomly initialized modules as $\Phi$ as plug-in for $\theta_{\text{pre}}$. For example, additive method Adapter-FT (Houlsby et al., 2019) inserts small and trainable fully connected networks $\Phi$ after transformer sub-layers in $\theta_{\text{pre}}$. The reparameterization-based method LoRA-FT (Hu et al., 2021) employs a low-rank matrix decomposition to parameterize the weight updates, and $\Phi$ indicates parameters for the low-rank matrices.
@@ 2.3. Success Metrics
Now we introduce concrete attack effectiveness metrics for different attacks and propose stealthiness metrics for victims to audit the pre-trained model.
@@ 4.2.1. Performance Comparison
In addition, we observe the advantage introduced by model initialization in Table 3 by comparing Benign with Basic and the more aggressive Accelerated. We set $E_{\text{pre}}=1$ as a safe choice for the accelerated version on all datasets. There is a clear trend that the Val-PPL is decreasing and the privacy risk is increasing from Benign to Basic to Accelerated. The Accelerated is indeed a more aggressive strategy that pushes the starting point to memorization-only stage.
@@ 4.2.4. Effectiveness under Defense
Figure 8. Duplication statistics and MIA effectiveness with Full-Ref under deduplication defense on $D_{\text{ft}}$ for PTB. The stealthiness metric $S_{\text{gap}}=\Delta_{\text{ppl}}$ is linearly scaled for clear visualization. We randomly subsample non-membership samples for keeping the same size as the deduplicated $D_{\text{ft}}$ in MIA evaluation.
@@
Figure 9. MIA effectiveness on PTB dataset with $D_{\text{aux}}^{\text{dedup}}$ and $D_{\text{aux}}^{\text{dup}}$ as auxiliary data for training $\theta_{\text{ref}}$ in Benign or $\theta_{\text{pre/ref}}$ in PreCurious, and $|D_{\text{dedup/dup}}|/|D_{\text{ft}}|=1$ denotes the default $D_{\text{aux}}$ w/o deduplication.
Under the representative defense strategy of weight decay, we show in Table 5 that PreCurious is robust on privacy risk amplification even with a high coefficient that exceeds the typical selection.
Under the strict defense of DP fine-tuning (Yu et al., 2021; Li et al., 2021), we show in Table 6 that PreCurious model increases the AUC compared to the Benign model but has a smaller TPR@0.01FPR and better utility due to the warming-up. The overall risk compared to non-DP fine-tuning in Table 1 is significantly mitigated by DP, supported by more results w.r.t. various budgets in the Appendix Table 8.
@@ 4.2.6. Breaking-up the trade-off
In addition, we address the privacy-utility trade-off issue in Table 6 with the lagging strategy as shown in Figure 10. Even when the worst-case privacy is bounded by a strict DP, we show that $\epsilon=1$ is still not a perfect protection. This success is due to more iterations for achieving the same utility, and also because the larger gradient norm derived from PreCurious-lagging fully exploits the per-sample sensitivity to reflect the influence of each sample.
@@ Appendix A Experimental Setup and More Results
Our experiments were conducted on an Ubuntu 20.04.6 system with 8 NVIDIA Quadro RTX 8000 GPUs. The source code and other artifacts have been made available  [^2] https://github.com/Emory-AIMS/PreCurious.